# Unstable Unlearning: The Hidden Risk of Concept Resurgence in Diffusion Models

### v2

@@ 1 Introduction
Our work begins with a surprising observation: fine-tuning a diffusion model can re-introduce previously erased concepts (see Figure 1 for a striking yet representative example). This can occur even when fine-tuning is performed on seemingly unrelated concepts. This hidden vulnerability, which we call concept resurgence, poses a challenge to the current paradigm of composing model updates via incremental fine-tuning. In particular, while the current state of the art in concept unlearning may initially suppress the generation of unwanted concepts (e.g., harmful, biased or copyrighted images), a developer cannot presently guarantee that concept unlearning will prevent the accidental reintroduction of these concepts in later updates to the model. As a consequence, consumers who fine-tune a “safe” model might inadvertently reintroduce undesirable behavior.
@@
  • Demonstrating the prevalence of concept resurgence. Through a series of systematic experiments, we investigate the conditions under which concept resurgence occurs. We show that concept resurgence does not require fine-tuning on data which is similar to the unlearned concept(s), or that the fine-tuning set is chosen adversarially to “jailbreak" the model. Instead, we show that concept resurgence can occur under common and benign usage patterns. Even well-meaning engineers may unintentionally expose users to unsafe or unwanted content that was previously removed. Figure 1 presents a representative example of this phenomenon.
@@
  • Investigating the cause(s) of concept resurgence. Finally, we develop a simple toy model to facilitate a systematic investigation into why concept resurgence occurs. This model highlights the fundamental challenges of both detecting and avoiding concept resurgence, and suggests a number of promising avenues for future research.
Organization of the paper. Section 2 covers background and related work. In Section 3, we quantify the extent of concept resurgence across a variety of domains. In Section 4, we explore some of the factors that influence the severity of concept resurgence. Finally, in Section 5 we construct a stylized model to systematically investigate the fundamental drivers of concept resurgence.
@@ 2 Background and related work
We investigate four recently proposed unlearning algorithms: ESD [13], SDD [22], UCE [13], and MACE [27]. At a high level, the first two methods focus on fine-tuning either the cross-attention weights or all of the model parameters such that encountering the concept of interest results in “unconditional" sampling (i.e., sampling which is not conditioned on the unwanted prompt). The latter two used closed-form edits to modify the cross-attention weights – and MACE additionally fine-tunes the remaining model parameters – to erase the concept of interest. We discuss these algorithms in additional detail in Section 4.2.
@@ 3 Composing Updates Causes Concept Resurgence
As discussed in Section 1, the scale of modern diffusion models has motivated a new paradigm in which updates to pretrained models are incrementally composed to avoid retraining models from scratch. These updates broadly take the form of one of two interventions: either the model is updated to learn a new concept, or it is updated to “unlearn" an unwanted concept. The standard procedure for learning new concepts is to curate a dataset of images representing the new concept of interest and fine-tune the model on this dataset. Similarly, to unlearn an unwanted concept(s), an “unlearning" algorithm will typically update the weights of the pretrained model in an attempt to ensure that the model no longer generates content associated with that concept. These two steps may be repeatedly composed over the lifetime of a deployed model. This paradigm raises an important question:
@@
Our investigation into this question begins with four of the most recent and performant unlearning methods discussed in Section 2: MACE, UCE, SDD, and ESD. We apply these unlearning algorithms to four different concept unlearning tasks (celebrity erasure, copyright erasure, unsafe content erasure, and object erasure) and two different diffusion models (Stable Diffusion v1.4 and Stable Diffusion v2.1). We describe these tasks in detail below. For each task, we first apply one of the unlearning algorithms to erase the concept of interest, and then subsequently fine-tune the model on a random set of in-domain concepts. For example, in the context of celebrity erasure — where the goal of the erasure task is to “unlearn" the ability to generate images of a particular celebrity — we further fine-tune the resulting model on a random set of celebrity images (which exclude the unlearned celebrity). This simulates the real world paradigm of composing unlearning with unrelated fine-tuning steps, the latter of which are intended to help the model learn new concepts or otherwise improve performance. In particular, we do not fine-tune the model on adversarially chosen concepts, as our goal is to understand whether benign updates can degrade or otherwise alter performance. For work on adversarial attacks and/or jailbreaking of text-to-image diffusion models, see [29, 46, 10]. Additionally, we focus on settings where the models retained high utility after unlearning.
@@
Figure 2: Selected images generated by SD v1.4 after initially applying each unlearning algorithm (top row) and after subsequent fine-tuning (bottom row) in the celebrity unlearning task. In each case, the model initially unlearns the target concept; e.g., how to generate images of Andrew Garfield. However, fine-tuning on unrelated images can inadvertently reintroduce the erased concepts. We note that UCE is more robust to this phenomenon then the other three algorithms. We discuss this result in Section 4.2.
Celebrity erasure. Following [27], the first benchmark we consider is inducing the model to forget certain celebrities (the “erase set") while retaining the ability to generate others (the “retain set"). We benchmark Stable Diffusion v1.4 and v2.1 in combination with each unlearning algorithm on the task of unlearning $100$ celebrities, and then evaluate whether the model succeeds in generating images of these celebrities (e.g., after being prompted with “A portrait of [erased celebrity name]"). To ensure consistency, both the subtasks and prompts are identical to those in [27]; the full set of celebrities in each subtask, along with the prompts used to evaluate the model, are provided in Appendix C. We quantify model performance across three random seeds by separately computing the mean top-1 accuracy of the Giphy Celebrity Detector (GCD) [18] on both erased and retained celebrities. [^1] The GCD is a popular open source model for classifying celebrity images; [27] document that the GCD achieves $>99\%$ top-1 accuracy on celebrity images sampled from Stable Diffusion v1.4.
Table 1: Unlearning performance before and after fine-tuning for Stable Diffusion v1.4 (Part 1). Each metric is task-specific, and evaluates the ability to generate the unwanted concept (lower is better; see Section 3 for details).
Celebrity Copyright
Method Before FT After FT Before FT After FT
ESD 0.144 $\pm$ 0.011 0.950 $\pm$ 0.007 $0$$\pm$ 0.000 0.100 $\pm$ 0.067
@@
SDD 0.556 $\pm$ 0.203 0.965 $\pm$ 0.008 $0$$\pm$ 0.000 0.100 $\pm$ 0.033
UCE 0.001 $\pm$ 0.001 0.004 $\pm$ 0.002 $0$$\pm$ 0.000 $0$$\pm$ 0.000
Table 2: Unlearning performance before and after fine-tuning for Stable Diffusion v1.4 (Part 2). Each metric is task-specific, and evaluates the ability to generate the unwanted concept (lower is better; see Section 3 for details). Results for SDD on unsafe content are excluded as first-stage unlearning compromises the model’s ability to generate any images, including retained concepts.
Object Unsafe
Method Before FT After FT Before FT After FT
@@
Copyright erasure. Motivated by recent, well-publicized concerns regarding the ability of diffusion models to generate copyrighted content  [40, 41, 44, 49], the second task we consider is one in which we induce the model to unlearn a popular fictional character while retaining the ability to generate other characters. Specifically, we apply each of the four unlearning algorithms to Stable Diffusion v1.4 and v2.1 to unlearn the concept “Iron Man", and then evaluate whether subsequent fine-tuning reintroduces the ability to generate this character (e.g., after being prompted with “a pose of Iron Man in action."). The full set of retained characters and the prompts used to evaluate the model are provided in Appendix C. We quantify the model performance by prompting Molmo 7B-D [8], an open-source multimodal LLM, with the generated image and two questions: “Is [copyrighted character] in this image? Answer Yes or No." and “Who is in this image? State their name only.". We categorize the image as including the character if the response to the first prompt is “Yes" or the character name is correct. We perform this evaluation across three random seeds on the set of evaluation prompts.
@@
Object erasure. Finally, following [27], the final benchmark we consider is inducing the model to forget how to generate certain types of objects from the CIFAR10 dataset (the “erase set") while retaining the ability to generate others (the “retain set"). We apply each unlearning algorithm to Stable Diffusion v1.4 to erase three objects (automobiles, ships, and birds) simultaneously. We then evaluate whether the model can generate images of these objects and their synonyms (e.g., after being prompted with “a photo of the [erased object]"). Both the full set of erased objects and retained objects, along with the prompts used to evaluate the model, are provided in Appendix C. As in the celebrity erasure task, we adopt the set of concepts to be erased, evaluation prompts and other hyperparameters from [27]. [^2] The only exception is the Erase 5 Objects task, which we add to evaluate simultaneous erasure of multiple concepts. We quantify model performance by computing the CLIP accuracy across three random seeds on the set of evaluation prompts.
Evaluating concept resurgence. In each of these settings, we are primarily concerned with whether concept resurgence occurs, and, if it does, the rate at which it does so. We curate specific examples to characterize the severity of concept resurgence in Figure 2. We show concept resurgence can occur in striking and seemingly unpredictable ways across all four algorithms, running the risk that developers or users can inadvertently reintroduce harmful or unwanted content.
@@
In Table 1 and 2, we quantify the degree of resurgence across all four tasks and unlearning algorithms using the metrics described above. The degree of resurgence varies across the algorithms and tasks. ESD and SDD exhibit a large degree of concept resurgence across all tasks; in some cases benign fine-tuning reverses unlearning almost completely. For MACE we see a modest degree of concept resurgence across all four tasks, and for UCE we see a small amount of resurgence in the celebrity and object erasure tasks. These findings illustrate that concept resurgence occurs with striking regularity across both algorithms and domains. We emphasize that in many contexts, even rare concept resurgence presents unacceptable risks. In the remainder of this work, we characterize the factors that affect the severity of concept resurgence and investigate the root causes of this phenomenon.
4 Factors Influencing Concept Resurgence Severity
4.1 Scaling Unlearning Algorithms
A key desideratum for any unlearning algorithm is the ability to scale: ideally, the user can erase many concepts without retraining the model from scratch. All four unlearning algorithms we consider report the ability to simultaneously unlearn many concepts while maintaining utility on unrelated concepts. We analyze whether increasing the number of concepts which are unlearned leaves the resulting model more susceptible to concept resurgence. For the celebrity erasure task, we define four subtasks: erasing $1$, $5$, $10$, and $100$ celebrities. For the object erasure task, we define three subtasks: erase ship, erase three objects (automobile, ship, bird), and erase five objects (automobile, ship, bird, cat, and truck). We follow the same evaluation setup as described in Section 3 for both tasks. We omit the copyright task from this analysis because we found that the models were unable to unlearn more than one character without dramatically degrading performance on retained characters. [^3] In this case, we interpret the algorithm as having failed in the first unlearning step, and thus there is no potential resurgence to evaluate. Without this requirement, a model which simply outputs random noise would suffice to achieve perfect performance on any unlearning task. We also omit the unsafe content task, as it cannot be cleanly decomposed into discrete “subtasks" (e.g., individual celebrities, objects or characters).
The impact of increasing the number of unlearned concepts varies amongst the four algorithms. For ESD, there is clear increase in resurgence as the number of concepts unlearned increases (Figure 3). In contrast, for MACE, UCE, and SDD the level of resurgence was not impacted as the number of concepts increased (see Appendix E). We discuss the possible mechanisms at play in the following section.

### v3

@@ 1 Introduction
Our work begins with a surprising observation: fine-tuning a diffusion model can re-introduce previously erased concepts (see Figure˜1 for a striking yet representative example). This can occur even when fine-tuning is performed on seemingly unrelated concepts and when users prompt the model to generate a completely unrelated concept. This hidden vulnerability, which we call concept resurgence, poses a challenge to the current paradigm of composing model updates via incremental fine-tuning. In particular, while the current state of the art in concept unlearning may initially suppress the generation of unwanted concepts (e.g., harmful, biased or copyrighted images), a developer cannot presently guarantee that concept unlearning will prevent the accidental reintroduction of these concepts in later updates to the model. As a consequence, consumers who fine-tune a “safe” model might inadvertently reintroduce undesirable behavior.
@@
  • Demonstrating the prevalence of concept resurgence. Through a series of systematic experiments, we investigate the conditions under which concept resurgence occurs. We show that concept resurgence does not require fine-tuning on data which is similar to the unlearned concept(s), or that the fine-tuning set is chosen adversarially to “jailbreak" the model. Instead, we show that concept resurgence can occur under common and benign usage patterns. Even well-meaning engineers may unintentionally expose users to unsafe or unwanted content that was previously removed. Figure˜1 presents a representative example of this phenomenon.
@@
  • Investigating the cause(s) of concept resurgence. We analyze a linear score-based diffusion model to understand, in a provable setting, why concept resurgence occurs after unlearning. Our analysis identifies two key factors that govern the strength of resurgence during fine-tuning: (1) the projection overlap between the forgotten subspace and the gradient directions introduced during fine-tuning, and (2) a curvature-limited sensitivity bound that quantifies how small gradient components in low-curvature subspaces can induce disproportionately large parameter updates. Crucially, our results show that some degree of resurgence is inevitable whenever there is nonzero overlap between the fine-tuning gradient subspace and the forgotten subspace, even if the overlap is small. Moreover, resurgence is most pronounced at early diffusion steps where gradients are strongest, but can also be amplified at intermediate-to-late steps when curvature is low and residual alignment persists.
Organization of the paper. Section˜2 covers background and related work. In Section˜3, we quantify the extent of concept resurgence across a variety of domains. In Section˜4, we explore some of the factors that influence the severity of concept resurgence. Finally, in Section˜5 we construct a stylized model to provably investigate the fundamental drivers of concept resurgence.
@@ 2 Background and related work
We investigate seven recently proposed unlearning algorithms: ESD [13], SDD [22], UCE [13], MACE [27], SalUn [11], SHS [44], and EraseDiff [45]. At a high level, ESD and SDD focus on fine-tuning either the cross-attention weights or all of the model parameters such that encountering the concept of interest results in “unconditional" sampling (i.e., sampling which is not conditioned on the unwanted prompt). EraseDiff performs unlearning similarly via a bi-level optimization problem. MACE and UCE used closed-form edits to modify the cross-attention weights – and MACE additionally fine-tunes the remaining model parameters – to erase the concept of interest. SalUn and SHS both start by identify the most influential parameters related to the concepts being unlearned and then finetune those parameters. We discuss these algorithms in additional detail in Section˜4.2.
@@ 3 Composing Updates Causes Concept Resurgence
As discussed in Section˜1, the scale of modern diffusion models has motivated a new paradigm in which updates to pretrained models are incrementally composed to avoid retraining models from scratch. These updates broadly take the form of one of two interventions: either the model is updated to learn a new concept, or it is updated to “unlearn" an unwanted concept. The standard procedure for learning new concepts is to curate a dataset of images representing the new concept of interest and fine-tune the model on this dataset. Similarly, to unlearn an unwanted concept(s), an “unlearning" algorithm will typically update the weights of the pretrained model in an attempt to ensure that the model no longer generates content associated with that concept. These two steps may be repeatedly composed over the lifetime of a deployed model. This paradigm raises an important question:
@@
Our investigation into this question begins with seven of the most recent and performant unlearning methods discussed in Section˜2: MACE, UCE, SDD, ESD, SalUn, SHS, and EraseDiff. We apply these unlearning algorithms to four different concept unlearning tasks (celebrity erasure, copyright erasure, unsafe content erasure, and object erasure) and two different diffusion models (Stable Diffusion v1.4 and Stable Diffusion v2.1). We describe these tasks in detail below. For each task, we first apply one of the unlearning algorithms to erase the concept of interest, and then subsequently fine-tune the model on a random set of in-domain concepts. For example, in the context of celebrity erasure — where the goal of the erasure task is to “unlearn" the ability to generate images of a particular celebrity — we further fine-tune the resulting model on a random set of celebrity images (which exclude the unlearned celebrity). This simulates the real world paradigm of composing unlearning with unrelated fine-tuning steps, the latter of which are intended to help the model learn new concepts or otherwise improve performance. In particular, we do not fine-tune the model on adversarially chosen concepts, as our goal is to understand whether benign updates can degrade or otherwise alter performance. For work on adversarial attacks and/or jailbreaking of text-to-image diffusion models, see [29, 47, 9]. Additionally, we focus on settings where the models retained high utility after unlearning. We describe the fine-tuning datasets and training details in Appendix˜C.
@@
Celebrity erasure. Following [27], the first benchmark we consider is inducing the model to forget certain celebrities (the “erase set") while retaining the ability to generate others (the “retain set"). We benchmark Stable Diffusion v1.4 and v2.1 in combination with each unlearning algorithm on the task of unlearning $100$ celebrities, and then evaluate whether the model succeeds in generating images of these celebrities (e.g., after being prompted with “A portrait of [erased celebrity name]"). To ensure consistency, both the subtasks and prompts are identical to those in [27]; the full set of celebrities in each subtask, along with the prompts used to evaluate the model, are provided in Appendix˜C. We quantify model performance across three random seeds by separately computing the mean top-1 accuracy of the Giphy Celebrity Detector (GCD) [18] on both erased and retained celebrities. [^1] The GCD is a popular open source model for classifying celebrity images; [27] document that the GCD achieves $99\%$ top-1 accuracy on celebrity images sampled from Stable Diffusion v1.4.
Figure 2: Selected images generated by SD v1.4 after initially applying each unlearning algorithm (top row) and after subsequent fine-tuning (bottom row) in the celebrity unlearning task. In each case, the model initially unlearns the target concept, e.g., how to generate images of Andrew Garfield. However, fine-tuning on unrelated images can inadvertently reintroduce the erased concepts. We note that UCE is more robust to this phenomenon than the other three algorithms. We discuss this result in Section˜4.2 and provide examples for SHS, SalUn, and EraseDiff in Appendix˜A.
Method Celebrity Copyright
Before FT After FT Before FT After FT
ESD 0.144 $\pm$ 0.011 0.950 $\pm$ 0.007 0.000 $\pm$ 0.000 0.100 $\pm$ 0.067
@@
SDD 0.556 $\pm$ 0.203 0.965 $\pm$ 0.008 0.000 $\pm$ 0.000 0.100 $\pm$ 0.033
UCE 0.001 $\pm$ 0.001 0.004 $\pm$ 0.002 0.000 $\pm$ 0.000 0.000 $\pm$ 0.000
EraseDiff 0.000 $\pm$ 0.000 0.693 $\pm$ 0.019 0.000 $\pm$ 0.000 0.367 $\pm$ 0.033
SHS 0.075 $\pm$ 0.019 0.893 $\pm$ 0.054 0.000 $\pm$ 0.000 0.133 $\pm$ 0.033
SalUn 0.363 $\pm$ 0.082 0.939 $\pm$ 0.056 0.000 $\pm$ 0.033 0.100 $\pm$ 0.067
(a) Celebrity and Copyright Tasks
Method Object Unsafe
Before FT After FT Before FT After FT
@@
EraseDiff 0.002 $\pm$ 0.002 0.995 $\pm$ 0.001 0.317 $\pm$ 0.181 0.876 $\pm$ 0.017
SHS 0.399 $\pm$ 0.274 0.999 $\pm$ 0.001 0.403 $\pm$ 0.058 0.848 $\pm$ 0.024
SalUn 0.831 $\pm$ 0.531 0.913 $\pm$ 0.065 0.840 $\pm$ 0.217 0.872 $\pm$ 0.008
(b) Object and Unsafe Tasks
Table 1: Unlearning performance before and after fine-tuning for Stable Diffusion v1.4. Each metric is task-specific and evaluates the ability to generate the unwanted concept (lower is better; see Section˜3 for details). Results for SDD on unsafe content are excluded as first-stage unlearning compromises the model’s ability to generate any images, including retained concepts.
Copyright erasure. Motivated by recent, well-publicized concerns regarding the ability of diffusion models to generate copyrighted content  [40, 41, 43, 50], the second task we consider is one in which we induce the model to unlearn a popular fictional character while retaining the ability to generate other characters. Specifically, we apply each of the seven unlearning algorithms to Stable Diffusion v1.4 and v2.1 to unlearn the concept “Iron Man", and then evaluate whether subsequent fine-tuning reintroduces the ability to generate this character (e.g., after being prompted with “a pose of Iron Man in action."). The full set of retained characters and the prompts used to evaluate the model are provided in Appendix˜C. We quantify the model performance by prompting Molmo 7B-D [7], an open-source multimodal LLM, with the generated image and two questions: “Is [copyrighted character] in this image? Answer Yes or No." and “Who is in this image? State their name only.". We categorize the image as including the character if the response to the first prompt is “Yes" or the character name is correct. We perform this evaluation across three random seeds on the set of evaluation prompts.
@@
Object erasure. Finally, following [27], the final benchmark we consider is inducing the model to forget how to generate certain types of objects from the CIFAR10 dataset (the “erase set") while retaining the ability to generate others (the “retain set"). We apply each unlearning algorithm to Stable Diffusion v1.4 to erase three objects (automobiles, ships, and birds) simultaneously. We then evaluate whether the model can generate images of these objects and their synonyms (e.g., after being prompted with “a photo of the [erased object]"). Both the full set of erased objects and retained objects, along with the prompts used to evaluate the model, are provided in Appendix˜C. As in the celebrity erasure task, we adopt the set of concepts to be erased, evaluation prompts and other hyperparameters from [27]. [^2] The only exception is the Erase 5 Objects task, which we add to evaluate simultaneous erasure of multiple concepts. We quantify model performance by computing the CLIP accuracy across three random seeds on the set of evaluation prompts.
Evaluating concept resurgence. In each of these settings, we are primarily concerned with whether concept resurgence occurs, and, if it does, the rate at which it does so. We curate specific examples to characterize the severity of concept resurgence in Figure˜2. We show concept resurgence can occur in striking and seemingly unpredictable ways across all seven algorithms, running the risk that developers or users can inadvertently reintroduce harmful or unwanted content.
In Table˜1, we quantify the degree of resurgence across all four tasks and unlearning algorithms using the metrics described above. The degree of resurgence varies across the algorithms and tasks. ESD, SDD, SalUn, SHS, and EraseDiff all exhibit a large degree of concept resurgence across all tasks; in some cases benign fine-tuning reverses unlearning almost completely. For MACE we see a modest degree of concept resurgence across all four tasks, and for UCE we see a small amount of resurgence in the celebrity and object erasure tasks. These findings illustrate that concept resurgence occurs with striking regularity across both algorithms and domains. We emphasize that in many contexts, even rare concept resurgence presents unacceptable risks. In the remainder of this work, we characterize the factors that affect the severity of concept resurgence and investigate the root causes of this phenomenon.
Incidental Concept Resurgence In conducting our object experiments, we uncover an even more concerning type of concept resurgence – the model will output an unlearned concept when prompted to generate an image of a retained concept. This means that a user can be prompting the model for an unrelated concept, and an unlearned concept is generated. We term this incidental concept resurgence. For example, when generating an image of an airplane that was retained, the model generates an image of an automobile that was unlearned before fine-tuning (example shown in Figure˜3 and in Appendix˜A). Furthermore, we calculate the percentage of prompts on which this phenomenon occurs across all seven algorithms for our erase-three and erase-five object tasks. We find that ESD, UCE, MACE, and SDD all share this vulnerability on at least one of the tasks. Meanwhile, SalUn, SHS, and EraseDiff appear robust (Tab.˜4).
(a) Stable Diffusion v1.4
(b) MACE
(c) Additional Fine-tuning
Figure 3: Images generated by the prompt “A photo of an airplane." Stable Diffusion v1.4 successfully generates this image (a), and Mass Concept Erasure (MACE) which unlearned {cat, truck, automobile, ship, bird}, partially generates this concept with the wing on the ground. However, subsequent fine-tuning on an unrelated set of randomly selected object images reintroduces the ability to generate the target concept when prompting with a completely unrelated concept ( (c).
4 Factors Influencing Concept Resurgence Severity
4.1 Scaling Unlearning Algorithms
@@
A key desideratum for any unlearning algorithm is the ability to scale: ideally, the user can erase many concepts without retraining the model from scratch. All seven unlearning algorithms we consider report the ability to simultaneously unlearn many concepts while maintaining utility on unrelated concepts. We analyze whether increasing the number of concepts unlearned leaves the resulting model more susceptible to concept resurgence. For the celebrity erasure task, we define four subtasks: erasing $1$, 55, 1010, and $100$ celebrities. For the object erasure task, we define three subtasks: erase ship, erase three objects (automobile, ship, bird), and erase five objects (automobile, ship, bird, cat, and truck). We follow the same evaluation setup as described in Section˜3 for both tasks. We omit the copyright task from this analysis because we found that the models were unable to unlearn more than one character without dramatically degrading performance on retained characters. [^3] In this case, we interpret the algorithm as having failed in the first unlearning step, and thus there is no potential resurgence to evaluate. Without this requirement, a model which simply outputs random noise would suffice to achieve perfect performance on any unlearning task. We also omit the unsafe content task, as it cannot be cleanly decomposed into discrete “subtasks" (e.g., individual celebrities, objects or characters). The impact of increasing the number of unlearned concepts is only noticeable for ESD. For ESD, there is clear increase in resurgence as the number of concepts unlearned increases (Figure˜4). In contrast, for the other six algorithms, the level of resurgence was not impacted as the number of concepts increased (see Appendix˜E).